OpenStudy (melissa07):

Brian claims that his dog can do single-digit multiplication. He shows the dog 10 cards with simple multiplication problems that are either correct or incorrect. The dog correctly identifies the correctness of the card 7 times out of 10. If the dog were simply guessing, how many would the dog be expected to get correct?

OpenStudy (anonymous):

x^2-2x=-1, 3x-x-x+a-a=5, (x^2-1)/(x+1), 2x-(x+x)

ganeshie8 (ganeshie8):

for each trial : P(Success) = 1/2 P(Fail) = 1/2

ganeshie8 (ganeshie8):

expected value = n*p = 10*1/2 = 5

OpenStudy (melissa07):

oh okay thank you !